OpenStudy (anonymous):

The Remainder Theorem states that: Select one: a. if a polynomial f(x) is added by x – k, then the remainder is r = f(k). b. if a polynomial f(x) is multiplied by x – k, then the remainder is r = f(k). c. if a polynomial f(x) is divided by x – k, then the remainder is r = f(k). d. none of the above

OpenStudy (astrophysics):

Hint: \[\frac{ p(x) }{ (x-a) } = q(x)\] with remainder r(x), where p(x) is a polynomial

OpenStudy (anonymous):

none of the above ?

OpenStudy (astrophysics):

Compare my answer with all of the above.

OpenStudy (anonymous):

ok thanks

OpenStudy (astrophysics):

Np

OpenStudy (anonymous):

ok when i solve by x can i use synthetic division ?